OpenStudy (anonymous):

Find the dimensions of the largest right cicular cone that can be inscribed in a sphere of the radius 12 inches.

OpenStudy (anonymous):

try it master

OpenStudy (anonymous):

??

mathslover (mathslover):

this also include calculus right?

OpenStudy (anonymous):

right

mathslover (mathslover):

ok first of all when will be the right circular cone will be largest when radius = 12 inches

OpenStudy (anonymous):

yes

mathslover (mathslover):

very confusing .. @Callisto

OpenStudy (anonymous):

yeah super haha

OpenStudy (anonymous):

http://www.themathpage.com/acalc/applied.htm check out problem no. 5 its similar but in it we want to find volume

mathslover (mathslover):

yes i also searched for it .but found only volume :(

OpenStudy (anonymous):

@mathslover i think in this case we'll take lateral area and total area LA= pi*r*l TA= LA + BA where LA= lateral area BA = base area TA= pi*r*l + pi*r^2 =pi*r(l+r) now we can max or min its dimensions easily

mathslover (mathslover):

Hmn @annas I am seriously diverged in a complicated question .. well it seems that you are going right but not sure

OpenStudy (anonymous):

ok...